# Near-Optimal Adaptive Transforms for LLM Quantization

### v1

@@ 2.2.2 Loss Minimization
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$
where $\odot$ represents elementwise multiplication. The lower bound of the trace term in the loss is
@@
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\operatorname{diag}\left(\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{1}\right\|^{2},\dots,\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{d}\right\|^{2}\right)\bm{T}^{-\top}\right)$
$=$ $\ \sum_{j=1}^{d}\sum_{k=1}^{d}\left({\mathbf{e}_{j}{T}^{-1}}\mathbf{e}_{k}\right)^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|{T}^{-1}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
@@
$\geq$ $\ d\left(d^{-1}\sum_{k=1}^{d}\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Jensen)
@@
$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (14)
$=$ $\ \sum_{k=1}^{d}\left(\left\|\bm{S}^{\frac{1}{2}}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}\left(\mathbf{e}_{j}^{\top}\bm{H}^{\top}\mathbf{e}_{k}\right)^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}d^{-1}\right)^{2}$
@@
2.2.3 Discussion
Under our smooth modeling in Theorem˜2, for any orthogonal $\bm{T}$ (including the identity $\bm{T}=\mathbf{I}$ and the Hadamard $\bm{T}=\bm{H}$), the minimization objective
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\left(\mathbb{E}_{\bm{\eta}}\mathrm{diag}\left(\bm{\eta}\right)^{2}\right)\bm{T}\bm{y}$ (15)
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{S}^{2}\right)\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right),$
so the orthogonal transforms will not be helpful for reducing the quantization error. Also note that the non-trivial choices $\bm{U}^{\prime}=\bm{H}$ and $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ are both essential for reducing the error by at most $d$ times, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario. Either trivial choices of $\bm{U}^{\prime}=\mathbf{I}$ or $\bm{S}^{\prime}=\mathbf{I}$ will lead to the same suboptimal trace term $\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$, which is the same as the case of orthogonal $\bm{T}$.
2.3 Proof for Integer (INT) Data Types
2.3.1 Quantization Error Modeling
We model the quantization error using pseudo noise similar to DiffQ (defossez2022differentiable).
Theorem 3 (INT Quantization Error Modeling)
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
Proof Define the scalar quantizer of $b$-bit symmetric integer as $q\left(x\right)=\left(\left\lfloor xs^{-1}\right\rfloor+2^{-1}\right)s$, where $x\in\mathbb{R}$, scale $s\in\mathbb{R}_{\neq 0}$, and $xs^{-1}\in\left[-2^{b-1},2^{b-1}\right)$. To simplify the analysis, we model the quantizer as $q\left(x\right)=\left(xs^{-1}+\xi\right)s=x+s\xi$ with random noise $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we use the AbsMax scale $s=2^{1-b}\left\|\bm{\alpha}\right\|_{\infty}$. Define $\eta=2^{1-b}\xi$, then we have $s\xi=\eta\left\|\bm{\alpha}\right\|_{\infty}$. The quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=2^{1-b}\mathbb{E}_{\xi}\xi=0$.
2.3.2 Loss Minimization
Using Theorem˜3, the minimization objective is
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\left\|\bm{T}\bm{y}\right\|_{\infty}\bm{\eta}\right\|^{2}$ (16)
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\bm{\eta}\bm{\eta}^{\top}\right)\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
@@ 2.3.2 Loss Minimization
We can obtain a lower bound for the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)$. Because $s_{1}^{2}\geq\dots\geq s_{d}^{2}$ and $s_{1}^{\prime-2}\leq\dots\leq s_{d}^{\prime-2}$, using von Neumann’s trace inequality, we obtain:
@@
Equality is attained when $\bm{R}=\mathbf{I}$.
Next, we obtain the lower and upper bounds for the term $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}_{\infty}$. Express the expectations as $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\mathbb{E}_{\bm{y}}\max_{k}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\geq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ (Jensen’s inequality) and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\sum_{k=1}^{d}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$. By the relationships among the mean, maximum, and summation values, we have:
@@
where $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ and $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\mathbf{e}_{k}^{\top}\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\mathbf{e}_{k}=\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\mathbf{e}_{k}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$.
@@
Proof We provide proofs for Gaussian and Laplacian distributions in Sections˜A.1 and A.2, respectively.
@@
When $\bm{U}^{\prime}=\bm{H}$, the marginal second moment becomes $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|=\left\|\bm{S}^{\prime}\bm{H}^{\top}\mathbf{e}_{k}\right\|=\sum_{j=1}^{d}s_{j}^{\prime 2}d^{-1}=d^{-1}\operatorname{tr}\left(\bm{S}^{\prime 2}\right)=d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$, which is equalized for all $k$, and $\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ is minimized to $d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$. Taken together, by setting $\bm{R}=\mathbf{I}$ and $\bm{U}^{\prime}=\bm{H}$, the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$ is bounded by $\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ within a gap of $d$ such that
By the Cauchy-Schwarz inequality,
$\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\geq\left(\operatorname{tr}\left(\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)^{\frac{1}{2}}\left(\bm{S}^{\prime 2}\right)^{\frac{1}{2}}\right)\right)^{2}=\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ (24)
and the equality is attained when $\bm{S}^{\prime}\propto\bm{S}^{\frac{1}{2}}$. Without loss of generality, we can choose $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ to minimize both the lower and upper bounds of the quantization loss.
2.3.3 Discussion
Consider the case of $\bm{T}$ being orthogonal. $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=d$ and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{2}\right)$. For any distribution $\mathcal{D}_{\mathrm{y}}$, the bounds are $\operatorname{tr}\left(\bm{S}^{2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq d\operatorname{tr}\left(\bm{S}^{2}\right)$. For the Hadamard $\bm{T}=\bm{H}$ and a tail-bounded $\mathcal{D}_{\mathrm{y}}$, $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)}\operatorname{tr}\left(\bm{S}^{2}\right)$ with high probability. Because $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$, an orthogonal $\bm{T}$ is suboptimal, and the error bounds of an orthogonal $\bm{T}$ can be at most $d$ times larger than those of the optimal $\bm{T}$, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario.
3 Experiments
3.1 Algorithm
1
Input: weights $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$, activations $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$, damping ratio $\lambda\in\mathbb{R}$
Output: WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and pre-quantized weight $\hat{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ for each block $i=1,\dots,d_{\mathrm{in}}/d$
2

### v2

@@ A.3 Equation Derivations
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\operatorname{diag}\left(\bm{\eta}\right)\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\operatorname{diag}\left(\bm{\eta}\right)\right)\bm{T}^{-\top}\right)$ (Equation˜36)
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\mathbb{E}_{\eta}\eta^{2}\right)\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right).$
Equation˜20.
@@
$=$ $\ \operatorname{tr}\biggl(\bm{T}^{-1}\operatorname{diag}\left(\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{1}\right\|^{2},\dots,\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{d}\right\|^{2}\right)$ $\!\!\!\!\!\bm{T}^{-\top}\biggr)$
$=$ $\ \sum_{j=1}^{d}\sum_{k=1}^{d}\left({\mathbf{e}_{j}\bm{T}^{-1}}\mathbf{e}_{k}\right)^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|\bm{T}^{-1}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (Equation˜16)
@@
$\geq$ $\ d\left(d^{-1}\sum_{k=1}^{d}\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Jensen’s inequality)
@@
$=$ $\ \sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (Equation˜40)
$=$ $\ \sum_{k=1}^{d}\left(\left\|\bm{S}^{\frac{1}{2}}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}\right)^{2}$ (plug in)
$=$ $\ \sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}\left(\mathbf{e}_{j}^{\top}\bm{H}^{\top}\mathbf{e}_{k}\right)^{2}\right)^{2}$
$=$ $\ \sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}d^{-1}\right)^{2}$
$=$ $\ d\left(d^{-1}\sum_{j=1}^{d}s_{j}\right)^{2}$
@@
Equation˜21.
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ (42)
$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}$ (Equation˜18)
$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\bm{y}^{\top}\bm{T}^{\top}\mathrm{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\bm{T}^{-1}\mathrm{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}$
$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\bm{y}^{\top}\bm{T}^{\top}\mathrm{diag}\left(\bm{\eta}\right)^{2}\bm{T}\bm{y}$ (orthogonality of $\bm{T}$)
$=$ $\ \mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\left(\mathbb{E}_{\bm{\eta}}\mathrm{diag}\left(\bm{\eta}\right)^{2}\right)\bm{T}\bm{y}$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\bm{T}\bm{y}$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{y}$ (orthogonality of $\bm{T}$)
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}$
[…]
In the case of $\bm{U}^{\prime}=\mathbf{I}$:
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (Equation˜16)
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\left(\bm{S}^{\prime 2}\odot\mathbf{I}\right)\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (plug in)
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{S}^{\prime 2}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{S}^{2}\right).$
In the case of $\bm{S}^{\prime}=\mathbf{I}$:
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (Equation˜16)
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{U}^{\prime\top}\left(\left(\bm{U}^{\prime}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{U}^{\prime}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (plug in)
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{U}^{\prime\top}\bm{U}^{\prime}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{S}^{2}\right).$
Equation˜24.
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ (45)
$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\left\|\bm{T}\bm{y}\right\|_{\infty}\bm{\eta}\right\|^{2}$ (Equation˜23)
$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\bm{\eta}^{\top}\bm{T}^{-\top}\bm{T}^{-1}\bm{\eta}$
$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\operatorname{tr}\left(\bm{\eta}^{\top}\bm{T}^{-\top}\bm{T}^{-1}\bm{\eta}\right)$
$=$ $\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\operatorname{tr}\left(\bm{T}^{-1}\bm{\eta}\bm{\eta}^{\top}\bm{T}^{-\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\bm{\eta}\bm{\eta}^{\top}\right)\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
@@
Equation˜25.
$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)$ (46)
$=$ $\ \operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)$ (Equation˜16
$=$ $\ \operatorname{tr}\left(\bm{S}\bm{U}^{\top}\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)$
Because $s_{1}^{2}\geq\dots\geq s_{d}^{2}$ and $s_{1}^{\prime-2}\leq\dots\leq s_{d}^{\prime-2}$, using von Neumann’s trace inequality,
@@
The equality can be attained by choosing $\bm{R}=\mathbf{I}$.
Express the expectations as
$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\mathbb{E}_{\bm{y}}\max_{k}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\geq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ (Jensen’s inequality) (48)
and
$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\sum_{k=1}^{d}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}.$ (49)
By the relationships among the mean, maximum, and summation values, we have:
@@
where, by Equation˜36,
$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ (51)
and
$\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\mathbf{e}_{k}^{\top}\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\mathbf{e}_{k}=\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\mathbf{e}_{k}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}.$ (52)
@@
Proof We provide proofs for the Gaussian and Laplacian distributions in Sections˜A.5.1 and A.5.2, respectively.
@@
When $\bm{U}^{\prime}=\bm{H}$, using Equations˜36 and 52, the marginal second moment in Equation˜55 becomes
$\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=$ $\ \left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}=\left\|\bm{S}^{\prime}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}=\sum_{j=1}^{d}s_{j}^{\prime 2}d^{-1}=d^{-1}\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ (56)
$=$ $\ d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2},$
which is equalized for all $k$, and $\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ is minimized to $d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$. Finally, combining Equations˜50, 55 and 56,